# One Framework to Rule Them All: Unifying Multimodal Tasks with LLM Neural-Tuning

### v2

@@ Abstract
Large-scale models have exhibited remarkable capabilities across diverse domains, including automated medical services and intelligent customer support. However, as most large models are trained on single-modality corpora, enabling them to effectively process and understand multimodal signals remains a significant challenge. Current research often focuses on designing task-specific or scenario-specific tuning strategies, which limits the scalability and versatility. To address this limitation, we propose a unified framework that concurrently handles multiple tasks and modalities. In this framework, all modalities and tasks are represented as unified tokens and trained using a single, consistent approach. To enable efficient multitask processing, we introduce a novel tuning strategy termed neural tuning, inspired by the concept of sparse distributed representation in the human brain, where only specific subsets of neurons are activated for each task. Furthermore, to advance research in multimodal and multitask learning, we present a new benchmark, MMUD, which includes samples annotated with multiple task labels spanning reasoning segmentation, referring segmentation, image captioning, and text-to-image generation. By applying neural tuning to pretrained large models on the MMUD benchmark, we demonstrate the ability to handle multiple tasks simultaneously in a streamlined and efficient manner. All models, code, and datasets will be released publicly upon publication, fostering further research and innovation in this field.
@@ 1 Introduction
In revisiting the human cognitive process, which serves as the inspiration for artificial intelligence, we observe that humans inherently excel at multitask learning and effortlessly adapt to new tasks. A key factor underpinning this ability is the principle of Sparse Distributed Representation (SDR), or the Sparse Coding Hypothesis (SCH), which posits that information is represented in a way where only a small fraction of neurons are active at any given time [4, 5]. This principle is believed to enhance the efficiency of information processing and enable the brain to form unique and robust representations of complex data. Motivated by these findings in neuroscience, we propose a unified framework for multitask and multimodal learning in large language models (LLMs), as illustrated in Figure 1. Our framework introduces two key innovations:
Unified Tokenization for Multimodal Multitask Learning. We formulate all tasks, as well as their multimodal data inputs and outputs, into a unified token-based representation. Unlike previous approaches that rely on cross-attention mechanisms to facilitate multimodal interactions [6, 7], our framework directly feeds the tokenized inputs into the model, leveraging the pretrained model’s self-attention mechanism [8]. This allows the model to compute relationships across and within modalities in a holistic manner, enhancing its ability to understand inter- and intra-modal dependencies. By unifying tasks and modalities in this manner, we train all multimodal tasks simultaneously, enabling the model to capture shared patterns across different input data. This approach not only simplifies the tuning process but also mitigates task-specific biases, as the model learns to generalize from diverse input-output mappings. Consequently, our method allows pretrained models to handle multiple tasks with a single causal language modeling objective, significantly reducing computational overhead compared to prior methods.
Neural Tuning Strategy. To further improve the efficiency and adaptability of multimodal multitask learning, we introduce neural tuning, a novel tuning strategy inspired by SDR. In this strategy, only a subset of neurons is activated for each task, mimicking the sparse activation patterns observed in the human brain. Neurons activated by different tasks consist of two components: shared neurons, which capture common features across tasks, and task-specific neurons, which encode unique characteristics for individual tasks. This design not only enhances multitask learning efficiency but also reduces computational costs by selectively activating neurons. Furthermore, by avoiding additional cross-attention mechanisms and complex decoders, our framework ensures simplicity and computational efficiency while maintaining strong performance across tasks.
@@
Figure 1: The overview of our proposed multitask and multimodal tuning framework. All inputs and outputs are token-based, encompassing both texts and images. The model generates specific tokens for different tasks, such as $<$OBJ$>$ for segmentation and $<$GEN$>$ for text-to-image generation. During tuning, a new sparse task network is introduced to emulate SDR and provide task guidance for pretrained LLMs. The entire LLM remains frozen, with only the newly introduced parameters being tunable.
@@ 2.1 Multimodal Tuning for Large Models
Recently, large models have demonstrated remarkable potential across diverse domains, not only in practical applications but also as tools for advancing scientific inquiry. However, training such models directly on multimodal data is often infeasible due to significant hardware and data constraints. To address this, various multimodal tuning strategies have emerged, primarily leveraging pre-trained language models to endow them with multimodal perception capabilities. The core concept of these approaches is to preserve the original model parameters by keeping them frozen, while adapting newly introduced components through lightweight methods such as adapters or specialized architectures. For example, Flamingo [9] retains a frozen image encoder and incorporates multimodal information via gated cross-attention, showcasing the ability of large language models (LLMs) to process multiple modalities. BLIP-2 [10], on the other hand, introduces the Q-Former module, a sophisticated mechanism designed to align image and text embeddings by extracting shared semantic representations, which has proven highly effective and widely adopted. Similarly, FROMAGe [11] employs linear transformations to bridge textual and visual features, enabling seamless translation between these domains. More recently, vision-language LLMs (LVLMs) such as LLaVA [12] and Qwen-VL [13] have demonstrated exceptional multimodal reasoning and understanding capabilities. Our proposed framework and the novel neural tuning strategy build on the foundations of multimodal tuning for LLMs. However, our approach emphasizes multitask learning in a more elegant and concise manner, particularly excelling in reasoning and segmentation tasks under complex scenarios.
@@ 2.2 Referring Segmentation and Reasoning Segmentation
Among multimodal understanding tasks, referring segmentation is a foundational challenge that involves segmenting objects in an image based on textual instructions. This task evaluates a model’s ability to establish connections between fine-grained visual details and linguistic cues, making it a crucial benchmark for assessing multimodal understanding. Earlier approaches typically utilized a text encoder alongside a U-Net-like [14] vision backbone to produce segmentation masks, as exemplified by methods such as LAVT [15] and SLViT [6]. With the advent of large models, however, referring segmentation has become relatively trivial, as simple textual instructions fail to challenge the advanced reasoning capabilities of these models. To address this limitation, the concept of complex reasoning segmentation was introduced. This paradigm extends beyond basic segmentation by incorporating complex questions about image content that require sophisticated reasoning to answer. The output is not limited to segmentation masks but also includes detailed textual responses to these complex queries. Furthermore, the model must identify and segment the objects relevant to its answers. For example, LISA [16] introduced a novel architecture with specialized tokens for segmentation, while PixelLM [2] proposed a newly designed codebook to support multi-instance segmentation. In our work, reasoning segmentation is a pivotal task for evaluating the multimodal reasoning and processing capabilities of our framework.
@@ 2.3 Text-to-Image Synthesis
For text-to-image synthesis, there are generally two approaches highly related to our work: vector quantized generative adversarial network (VQGAN) related methods [17] and diffusion-based methods [3, 18]. VQGAN aims to map images into a discrete latent space while diffusion models simulates a diffusion process, where data is gradually transformed from a simple prior distribution (like Gaussian noise) to the complex target distribution. State-of-the-art text-to-image models such as GLIDE [19], DALL-E [20], and Imagen [21] leverage advanced techniques in VQGAN and diffusion models to generate high-quality, diverse images. In our work, we primarily employ VQGAN for synthesizing images, but we also explore the potential of combining our approach with diffusion networks for image generation.
@@ 3 Method
In this section, we first describe the detailed design of our framework with neural tuning, and then outline the specifics of multitask training.
@@ 3.1 All-in-token Multimodal Paradigm
The overall pipeline of our proposed neural tuning framework is illustrated in Figure 1. The inputs from the different modalities are tokenized and processed concurrently, emulating the ability of humans to seamlessly handle multiple tasks. We also introduce a novel sparse task network into pretrained models, which functions as either the Sparse Distributed Representation (SDR) or the Sparse Coding Hypothesis (SCH), through which only a subset of neurons are activated, tailored to the specific task at hand. This paradigm enables the model to take both image and text inputs, generating task-specific outputs such as multi-instance referring segmentation ($<$OBJ$>$) and image generation ($<$GEN$>$).
@@
Specifically, for a multimodal input consisting of images and sentences, we first embed the text into $I_{txt}=\{I_{txt}^{l}\}_{l=1}^{L_{t}}$, where $L_{t}$ denotes the length of the text. Next, we use a pretrained vision encoder to extract features from the image, resulting in $I_{img}=\{I_{img}^{l}\}_{l=1}^{L_{i}}$ where $L_{i}$ representing the number of image patches. To integrate the multimodal input, we concatenate the text embeddings and image features (when the task requires visual input) to form the final input for the pretrained large language model (LLM): $I=[I_{img};I_{txt}]$. In this scheme, the textual and visual modalities interact through the LLM’s self-attention mechanism, where the output $O$ is computed as $O=SoftMax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ with the Query($Q$), Key($K$), and Value($V$) derived from the concatenated multimodal input $I$. Unlike prior works that employ cross-attention for modality interaction [6, 15], where $Q$, $K$, and $V$are derived from different modalities, our self-attention approach offers several advantages. Not only does it simplify the architecture by avoiding explicit cross-modal mappings, but it also enables the model to compute both inter- and intra-modal relationships. This dual interaction within the self-attention framework enhances the model’s ability to understand and fuse multimodal information more effectively.
@@
In the output stage of the LLM, we introduce new task-specific tokens alongside the original textual tokens to handle multimodal tasks. For example, we incorporate $<$OBJ$>$ tokens for segmentation tasks and $<$GLB$>$ tokens for text-to-image synthesis. These task-specific tokens are then fed into their respective decoders, enabling the model to generate the appropriate outputs for each task. This approach unifies the input and output format across all tasks into an all-in-token scheme, simplifying the integration of additional tasks and modalities. As a result, the system becomes more flexible and scalable, allowing for easy extension to new multimodal tasks without significant architectural changes.
@@ 3.2 Neural Tuning for Large Models
where $Z_{k-1}^{u}$ is the sparse vector from the last layer. This approach allows the sparse vector to not only receive information from the previous layer but also be aware of the current layer’s information. To enable sparse distributed representation during model flow, only a subset of the sparse vector’s nodes are activated for various tasks. We first randomly sample an activation rate $r$ from a normal distribution $p(r)=N(r;\beta,(0.1\beta)^{2})$, where $\beta\in(0,1)$ is a predefined hyperparameter [^1] More activation patterns are discussed in Section 5.5. Then, we activate corresponding neurons as follows:
@@
$Q^{\prime}$, $K$, and $V^{\prime}$ are then employed for the vanilla self-attention in pretrained LLMs. Overall, the pretrained LLM is responsible for the main inference, while the sparse task network handles the execution of different tasks. Interestingly, our experiments reveal that more related tasks tend to activate a greater number of the same neurons (e.g., vanilla referring segmentation and reasoning segmentation). This finding highlights the relationship between different tasks, as illustrated in Figure 2(b).
@@ 3.3 Multitask Training
As different tasks are unified into an all-in-one token manner, the tuning procedure can be conducted using a simple causal language modeling approach. Cross-entropy loss is employed as the loss function ($\mathcal{L}_{txt}$) for next-token prediction:
@@
For segmentation-related tasks, the embeddings corresponding to object tokens $<$OBJ$>$, $H_{seg}$, are collected and fed to a lightweight decoder for corresponding mask generation:
@@
where $W_{seg}$ represents the learnable parameters in the mask decoder. Different from previous methods [16], which employ pretrained SAM [22] for predicting masks, our designed decoder is much lighter. The mask decoder consists of three convolutional layers for upscaling, each followed by three self-attention layers for mask decoding. Since the models can generate multiple segmentation tokens at a time, we can easily perform multi-instance segmentation. Following previous works [16, 2], DICE loss [23] is employed in our framework to guide the segmentation tasks:
@@
where $\lambda_{seg}$ and $\lambda_{gen}$ are coefficients to balance the numerical scales of different losses.
Figure 3: Pipeline for MMUD Dataset Generation. First, GPT-4v is used to generate captions describing the image contents and reasoning question-answer pairs. Then to ensure dataset quality, meaningless cases are filtered out. Finally, $<$OBJ$>$ tokens are manually appended to objects in the answers to help large models better understand the relationships between images and text.

### v3

@@ Abstract
Large-scale models have exhibited remarkable capabilities across diverse domains, including automated medical services and intelligent customer support. However, as most large models are trained on single-modality corpora, enabling them to effectively process and understand multimodal signals remains a significant challenge. Current research often focuses on designing task-specific or scenario-specific tuning strategies, which limits the scalability and versatility. To address this limitation, we propose a unified framework that concurrently handles multiple tasks and modalities. In this framework, all modalities and tasks are represented as unified tokens and trained using a single, consistent approach. To enable efficient multitask processing, we introduce a novel tuning strategy termed neural tuning, inspired by the concept of sparse distributed representation in the human brain, where only specific subsets of neurons are activated for each task. Furthermore, to advance research in multimodal and multitask learning, we present a new benchmark, MMUD, which includes samples annotated with multiple task labels spanning reasoning segmentation, referring segmentation, image captioning, and text-to-image generation. By applying neural tuning to pretrained large models on the MMUD benchmark, we demonstrate the ability to handle multiple tasks simultaneously in a streamlined and efficient manner. All codes and datasets will be released at https://github.com/kiva12138/NeuralTuning.
@@
{graphicalabstract}[Uncaptioned image]
{highlights}
Propose a framework that unifies multimodal tasks with a concise all-in-token manner.
The framework enhances the flexibility and scalability of LLMs.
Introduce neural tuning, a efficient yet effective task-specific tuning strategy.
Present MMUD, a multimodal dataset for tasks like referring segmentation and generation.
Reach state-of-the-art in performance by involving neural tuning on MMUD.
@@ 1 Introduction
In revisiting the human cognitive process, which serves as the inspiration for artificial intelligence, we observe that humans inherently excel at multitask learning and effortlessly adapt to new tasks. A key factor underpinning this ability is the principle of Sparse Distributed Representation (SDR), or the Sparse Coding Hypothesis (SCH), which posits that information is represented in a way where only a small fraction of neurons are active at any given time [4, 5]. By activating only the necessary subset of neurons, SDR reduces energy consumption and enhances the brain’s capacity to form unique, robust representations of complex and diverse inputs. Motivated by these findings in neuroscience, we propose a unified framework for multitask and multimodal learning in large language models (LLMs), as illustrated in Figure 1. Our framework introduces two key innovations:
Unified Tokenization for Multimodal Multitask Learning. We formulate all tasks, as well as their multimodal data inputs and outputs, into a unified token-based representation. Unlike previous approaches that rely on cross-attention mechanisms to facilitate multimodal interactions [6], our framework directly feeds the tokenized inputs into the model, leveraging the pretrained model’s self-attention mechanism. This allows the model to compute relationships across and within modalities in a holistic manner, enhancing its ability to understand inter- and intra-modal dependencies. By unifying tasks and modalities in this manner, we train all multimodal tasks simultaneously, enabling the model to capture shared patterns across different input data. This approach not only simplifies the tuning process but also mitigates task-specific biases, as the model learns to generalize from diverse input-output mappings. Consequently, our method allows pretrained models to handle multiple tasks with a single causal language modeling objective, significantly reducing computational overhead compared to prior methods. Furthermore, by avoiding additional cross-attention mechanisms and complex decoders, our framework ensures simplicity and computational efficiency while maintaining strong performance across tasks.
Neural Tuning Strategy. To further improve the efficiency and adaptability of multimodal multitask learning, we introduce neural tuning, a novel tuning strategy inspired by SDR. In this strategy, only a subset of neurons is activated for each task, mimicking the sparse activation patterns observed in the human brain. Neurons activated by different tasks consist of two components: shared neurons, which capture common features across tasks, and task-specific neurons, which encode unique characteristics for individual tasks. This approach aligns with the intrinsic strengths of SDR, enabling efficient multitask learning by minimizing task interference and reusing shared features while reserving task-specific pathways. By emulating this biologically inspired mechanism, our framework enhances scalability and adaptability, making it particularly suitable for handling the complexity of modern multimodal and multitask challenges in LLMs.
@@
Figure 1: The overview of our proposed multitask and multimodal tuning framework. All inputs and outputs are token-based, encompassing both texts and images. The model generates specific tokens for different tasks, such as <<OBJ>> for segmentation and <<GEN>> for text-to-image generation. During tuning, a new sparse task network is introduced to emulate SDR and provide task guidance for pretrained LLMs. The entire LLM remains frozen, with only the newly introduced parameters being tunable.
@@ 2.1 Multimodal Tuning for Large Models
LLMs exhibit exceptional versatility across domains, but direct multimodal training is often limited by substantial hardware and data demands. To mitigate this, multimodal tuning strategies leverage pre-trained LLMs without altering their core parameters. For example, BLIP-2 [7] employs a Q-Former to align image and text embeddings, and FROMAGe [8] uses linear projections to bridge modalities. LVLMs like LLaVA [9] and Qwen-VL [10] enhance reasoning via cross-attention mechanisms, unifying input representations across modalities. However, these methods mainly focus on modality interaction at the input level. VisionLLM v2 [11] advances this by enabling unified processing across tasks, though it lacks task-specific fine-tuning strategies. In contrast, our framework introduces an “all-in-token” paradigm and a novel neural tuning approach tailored for multitask learning. This removes the need for explicit cross-modal mappings, simplifying the architecture while capturing both inter- and intra-modal relations through a unified self-attention mechanism. Unlike VisionLLM v2 [11], our method emphasizes multitask efficiency, enabling parameter-efficient tuning with strong performance across segmentation and reasoning tasks. While recent all-in-token approaches propose unified multimodal tokens, our key innovation lies in the neural tuning mechanism specifically designed for multitask adaptation—achieving scalability, efficiency, and improved accuracy across diverse tasks.
@@ 2.2 Referring Segmentation and Reasoning Segmentation
Referring segmentation, a key multimodal task, involves segmenting image regions based on textual instructions, testing a model’s ability to align fine-grained visual details with language. Earlier methods, such as LAVT [12] and SLViT [6], paired text encoders with U-Net-like [13] vision backbones to generate masks. However, with the rise of large models, simple referring instructions pose limited challenges. To address this, complex reasoning segmentation has emerged—requiring models to answer detailed image-related questions and provide corresponding segmentations. For instance, LISA [14] introduces segmentation-specific tokens, while PixelLM [2] uses a custom codebook for multi-instance segmentation. In our framework, reasoning segmentation serves as a core benchmark for evaluating multimodal reasoning and understanding.
@@ 2.3 Text-to-Image Synthesis
For text-to-image synthesis, there are generally two approaches highly related to our work: vector quantized generative adversarial network (VQGAN) related methods [15] and diffusion-based methods [3, 16]. VQGAN aims to map images into a discrete latent space while diffusion models simulates a diffusion process, where data is gradually transformed from a simple prior distribution (like Gaussian noise) to the complex target distribution. State-of-the-art text-to-image models such as DALL-E [17] leverage advanced techniques in VQGAN and diffusion models to generate high-quality, diverse images. In our work, we primarily employ VQGAN for synthesizing images, but we also explore the potential of combining our approach with diffusion networks for image generation.
@@ 3 Method
In this section, we first describe the detailed design of our unified framework with neural tuning, and then outline the specifics of multitask training.
@@ 3.1 All-in-token Multimodal Paradigm
The overall pipeline of our proposed all-in-token framework is illustrated in Figure1. To align with humans’ ability to handle multiple tasks and modalities concurrently, inputs from different modalities are tokenized and processed in parallel. Inspired by the Sparse Distributed Representation (SDR) or Sparse Coding Hypothesis (SCH) of the human brain—where only a subset of neurons is activated for a specific task—we also introduce a novel sparse task network into pretrained models (further discussed in Section 3.2). This paradigm enables the model to process both image and text inputs and generate task-specific outputs, such as multi-instance referring segmentation (<<OBJ>>) and image generation (<<GEN>>).
Specifically, for a multimodal input consisting of images and sentences, we first embed the text into $I_{txt}=\{I_{txt}^{l}\}_{l=1}^{L_{t}}$, where $L_{t}$ denotes the length of the text. For image inputs, we adapt to the pretrained image encoders by first dividing the image into patches, with each patch representing a small region of the image. We then use a frozen pretrained vision encoder (such as CLIP [18]) to extract features from these patches, resulting in $I_{img}=\{I_{img}^{l}\}_{l=1}^{L_{i}}$, where $L_{i}$, where $L_{i}$ denotes the number of image patches.
@@
We then use a pretrained vision encoder to extract features from the image, resulting in $I_{img}=\{I_{img}^{l}\}_{l=1}^{L_{i}}$, where $L_{i}$ represents the number of image patches. To integrate the multimodal input, we concatenate the text embeddings and image features (when visual input is required for the task) to form the final input for the pretrained large language model (LLM): $I=[I_{img};I_{txt}]$. In this scheme, the textual and visual modalities interact through the LLM’s self-attention mechanism, where the output $O$ is computed as $O=SoftMax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ with the Query($Q$), Key($K$), and Value($V$) derived from the concatenated multimodal input $I$.
Unlike prior works that use cross-attention for modality interaction [6, 12], where $Q$, $K$, and $V$ are derived from different modalities, our all-in-token approach offers several advantages. It not only simplifies the architecture by avoiding explicit cross-modal mappings, but also enables the model to compute both inter- and intra-modal relationships. This dual interaction within the self-attention framework enhances the model’s ability to effectively understand and fuse multimodal information. Furthermore, when incorporating additional modalities (such as audio), this approach eliminates the need for designing complex cross-attention schemas and allows for seamless extension to new modalities by simply concatenating the corresponding modality tokens.
@@
In the output stage, we introduce new task-specific tokens alongside the original textual tokens to handle multimodal tasks. For instance, we incorporate <<OBJ>> tokens for segmentation tasks and <<GLB>> tokens for text-to-image synthesis. These task-specific tokens are then passed to their respective decoders, enabling the model to generate appropriate outputs for each task. This approach unifies the input and output formats across all tasks into an all-in-token scheme, simplifying the integration of additional tasks and modalities. Consequently, the system becomes more flexible and scalable, allowing for seamless extension to new multimodal tasks without significant architectural modifications.
@@ 3.2 Neural Tuning for Large Models
where $Z_{k-1}^{u}$ is the sparse vector from the last layer. This approach allows the sparse vector to not only receive information from the previous layer but also be aware of the current layer’s information. To enable sparse distributed representation like human brain during model flow, only a subset of the sparse vector’s nodes are activated for various tasks. We first randomly sample an activation rate $r$ from a normal distribution $p(r)=N(r;\beta,(0.1\beta)^{2})$, where $\beta\in(0,1)$ is a predefined hyperparameter [^1] More activation patterns are discussed in Section 5.5. Then, we activate corresponding neurons as follows:
@@
$Q^{\prime}$, $K$, and $V^{\prime}$ are then employed for the vanilla self-attention in pretrained LLMs. Overall, the pretrained LLM is responsible for the main inference, while the sparse task network handles task-specific execution, just like different parts in human brains.
Building on the described activation mechanism, neurons activated by different tasks consist of two components: shared neurons, which capture common features across tasks, and task-specific neurons, which encode the unique characteristics of individual tasks. Interestingly, our experiments reveal that tasks with higher relatedness tend to activate a greater number of shared neurons $Z_{k}^{a}$. For example, tasks like vanilla referring segmentation and reasoning segmentation exhibit significant overlap in neuron activation patterns, as they share common underlying features, such as object detection and spatial reasoning. Figure 2(b) provides a detailed visualization of these activation patterns. However, as we could draw from the visualization, some neurons are rarely activated across different tasks. This phenomenon can be attributed to two factors. Firstly, certain neurons are highly specialized for tasks that are less emphasized or represented in current experimental setups, resulting in fewer activations. Secondly, sparse activation ensures that only the most relevant neurons are utilized for a given task. This approach not only enhances computational efficiency but also prevents over-reliance on any single component, enabling the model to maintain adaptability across diverse tasks.
This visualization not only supports the design rationale of our sparse task network but also offers insights into how the model dynamically allocates resources to shared and task-specific components. Such behavior aligns with the human brain’s efficiency in reusing common cognitive processes while adapting to task-specific demands.
@@ 3.3 Multitask Training
The whole pipeline for our proposed multitask training is shown in Figure 3. As different tasks are unified into an all-in-one token manner, the tuning procedure can be conducted using a simple causal language modeling approach. Cross-entropy loss is employed as the loss function ($\mathcal{L}_{txt}$) for next-token prediction:
@@
Figure 3: The pipeline of our proposed framework for multitask training. After processing multimodal tokens, the LLM generates corresponding task tokens along with respective explanations. For segmentation-related tasks, a lightweight three-layer decoder is used for mask generation, while for image synthesis tasks, VQGAN is employed to generate images in an autoregressive manner.
For segmentation-related tasks, while it is feasible to use the same architecture as the image-synthesis task, we prioritize efficiency by proposing a separate lightweight decoder for segmentation mask generation. Specifically, the embeddings corresponding to object tokens, <<OBJ>>, denoted as $H_{seg}$, are extracted and fed into the lightweight decoder:
@@
where $W_{seg}$ represents the learnable parameters in the mask decoder. Different from previous methods [14], which employ pretrained SAM [19] for predicting masks, our designed decoder is much lighter. The mask decoder consists of three convolutional layers for upscaling, each followed by three self-attention layers for mask decoding. Since the models can generate multiple segmentation tokens at a time, we can easily perform multi-instance segmentation. Following previous works [14, 2], DICE loss is employed in our framework to guide the segmentation tasks:
@@
where $\lambda_{seg}$ and $\lambda_{gen}$ are coefficients to balance the numerical scales of different losses. Although segmentation and generation tasks are the primary illustration for multitask scalability in our current approach, the method can be easily extended to additional tasks by introducing new corresponding task-specific tokens.
Figure 4: Pipeline for MMUD Dataset Generation. First, GPT-4v is used to generate captions describing the image contents and reasoning question-answer pairs. Then to ensure dataset quality, meaningless cases are filtered out. Finally, <<OBJ>> tokens are manually appended to objects in the answers to help large models better understand the relationships between images and text.